OpenStudy (anonymous):

P=(RT/V) + (a+bT)/(V^2) What is the partial derivative of ∂(V)/∂(T)?

OpenStudy (thomas5267):

I don't know whether this will be complicated by Maxwell's relation which I don't understand. If this can be treated as a normal function, I will express V in terms of T before differentiating.

OpenStudy (irishboy123):

might be hard getting that V isolated, so you can use the inverse of \[\left (\frac{∂T}{∂V} \ \right )_P\] holding P constant, as the T is easier to isolate; or you could even do an *implicit* \[\frac{∂}{∂T} \] on the whole shebang, again holding P constant (so ∂P/∂T = 0, ∂T/∂T = 1 etc) not really thought about which would be neater so i'd do both -- and make sure they gave the same answer.